# Adaptive Gradient Calibration for Single-Positive Multi-Label Learning in Remote Sensing Image Scene Classification

### v1

@@ IV-A2 AID-multilabel
It extends the AID scene classification dataset [47] with multi-label annotations. The dataset comprises 3,000 high-resolution RGB aerial image patches ($600\times 600$ pixels) across 17 classes, collected from China, U.S., and several European countries. We exclude the extremely rare “mobile home” class (only 2 samples), resulting in 16 classes. Following [13], 20% of the data is held out for testing. We then use 25% of the remaining training set for validation. Dataset statistics are provided in Table II. For simplicity, in the following we use AID to refer to AID-multilabel.
TABLE II: AID Multilabel Classes and Related Flip Rates $\beta_{c}$ on the
Training Subset in the Case of Random SPML Label Noise.
Macro and Micro Averages Are Also Provided, as Well
as the Average Number of True Positive
Labels per Image
Class name Random $\beta_{c}$ Support
Airplane 0.83 59
Bare-soil 0.80 883
Buildings 0.82 1313
Cars 0.84 1211
Chaparral 0.72 58
Court 0.85 207
Dock 0.91 163
Field 0.74 129
Grass 0.79 1366
Pavement 0.81 1408
Sand 0.47 153
Sea 0.77 133
Ship 0.86 176
Tanks 0.83 65
Trees 0.78 1441
Water 0.84 509
Micro average 0.81
Macro average $\bar{\beta}$ 0.79
Average #labels/image 5.2
IV-B Experimental Setup
We adopt ResNet-34 and ResNet-50 as backbone networks for the reBEN and AID datasets, respectively, with batch sizes of 512 and 32. We choose ResNet-34 for reBEN as it has fewer parameters and lower computational complexity, which significantly reduces the overall runtime of our experiments. Despite being lighter, it shares a similar architecture with ResNet-50 and generally achieves comparable performance. Moreover, since our primary goal is to evaluate the effectiveness of the proposed method relative to baseline settings, rather than to maximize absolute accuracy, ResNet-34 provides a suitable and efficient compromise. Given the AID dataset’s limited size and RGB data, we initialize the network with ImageNet-pretrained weights. In this setup, we consider the realistic scenario of absence of clean labels for training and hyperparameter tuning. Hence, all methods’ hyperparameters were tuned via grid search by evaluating their mAP score on the noisy validation set. To manage the computational burden associated with the large number of methods and hyperparameters, we performed this tuning process using a subset of the respective datasets (20% of training/validation sets for reBEN and 50% for AID) and for a reduced number of epochs (10 epochs instead of 70 as in the final runs). For AdaGC, we set the Mixup Beta distribution parameter $\alpha$ to 1, and the GC loss weight $\lambda$ to 3, 5, and 8 for reBEN-Random, reBEN-Dominant, and AID, respectively, selected via hyperparameters tuning.
For the final evaluation, we utilized the optimal hyperparameters identified during the tuning phase. Each model was then trained on the full training dataset for a total of 70 epochs, after which each model was evaluated on the test dataset. For the proposed AdaGC approach, the noisy validation mAP from the teacher model, as discussed in Section III-C, was used to detect the optimal epoch for initiating GC with pseudo-label generation. To ensure the robustness and statistical reliability of our results, each method underwent three Monte Carlo runs, allowing us to quantify the variability in performance due to random weight initialization and dataset shuffling. We report the average values of all metrics along with their standard deviations (computed using an unbiased estimator).
For comprehensive evaluation, in addition to mAP, we report two ranking-based metrics, i.e., coverage and ranking loss (Rankloss), and four accuracy-based metrics, i.e., overall accuracy (OA), mean F1 score (mF1), mean precision (mprecision), and mean recall (mrecall). The metrics are defined as follows.
  • mF1 is the average of F1-scores computed per class:
  • Baseline methods: We considered several baselines [12], such as AN, AN-LS, Weak (WAN), and EPR. All use a BCE loss at their core, while EPR uses only the BCE positive term and adds a regularization term, assuming the a-priori knowledge of the average number of positive labels per sample (see Tables I and II).
  • General label noise methods: From this category, we considered the ELR approach adapted as in [10], the LL family of approaches (i.e., LL-R, LL-Ct, and $L$-Cp) [16], and the BoostLU [19] in combination with $L$-Ct.
  • PUL methods: We considered a single SOTA method from this category, i.e., Dist-PU [39], adapted for SPML by simply applying it to each class independently.
  • SPML-specific methods: The main comparison methods we considered for SPML are ROLE [12], GR Loss [18], LAGC [17], and Mutual label enhancement for sIngle-positive Multi-label lEarning (MIME) [20].
  • Upper bound methods: For analytical comparison with GT training, we considered two cases: i) training with GT multi-labels using a BCE loss (GT w/ BCE), and ii) the with true Negatives (IUN) approach [12], where in addition to the single positive labels, all true negatives are provided for training. IUN allows us to assess the performance in a hypothetical scenario where all true negatives have been retrieved. False negatives are ignored during training, meaning that the model is still provided with a single positive label for each image.
IV-C Results
@@
mAP (%) $\uparrow$ Coverage $\downarrow$ Rankloss ($\times 10^{2}$) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
@@
Proposed AdaGC 68.18 (0.35) 4.513 (0.012) 4.073 (0.033) 92.52 (0.02) 64.65 (0.37) 70.36 (0.84) 63.37 (0.82)
@@
mAP (%) $\uparrow$ Coverage $\downarrow$ Rankloss ($\times 10^{2}$) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
@@
Proposed AdaGC 60.22 (0.24) 5.618 (0.004) 6.926 (0.016) 90.15 (0.05) 58.78 (0.44) 61.28 (0.98) 62.48 (0.62)
@@
This section presents a detailed analysis of the experimental results reported in Tables III, IV, and V, comparing the performance of AdaGC against various baselines, PUL, SPML-specific, and general label noise methods across the reBEN and AID datasets under different SPML noise types. Several general trends are observable across all experimental scenarios. Our proposed AdaGC consistently achieves the best mAP values across all three evaluated scenarios. Furthermore, it demonstrates SOTA SPML performance in terms of coverage, rankloss, and mF1, highlighting its overall effectiveness and robustness in handling SPML learning for RS imagery. The upper-bound method IUN, which leverages knowledge of true negatives while only observing single positives, consistently achieves a good mAP. However, due to its bias towards single positive labels, it typically exhibits high mprecision but comparatively lower mrecall, which consequently impacts its mF1. The AN and AN-LS baselines are particularly sensitive to the inherent bias introduced by single positive labels and the prevalence of false negatives. This sensitivity is reflected in their high mprecision but significantly low mrecall, leading to a generally low mAP across all datasets. In contrast, the WAN method demonstrates greater robustness compared to AN and AN-LS. It achieves a better balance between mprecision and mrecall, making it the best-performing baseline method in terms of mAP across the considered scenarios.
EPR employs a distinct strategy by training solely on single positive samples and incorporating a regularization term to enforce the prediction of an average number of positive labels. This approach consistently results in EPR being the model with the highest mrecall across all datasets and noise types. However, given its reliance only on positive supervision, its mprecision is generally lower compared to methods that attempt to identify true negatives. Additionally, the consistently high mrecall scores of EPR are attributable to its unique reliance on a-priori knowledge of the average number of positive labels per image, a piece of information not utilized by other methods under comparison. Next subsections discuss the results obtained on each specific dataset and scenario.
@@
TABLE V: Test Performance Comparison of Different Methods on the AID Dataset. The Best Average Metric Values Are Reported in Bold, Second Bests Are in Italic. The Related Standard Deviations Are Reported in Brackets
mAP (%) $\uparrow$ Coverage $\downarrow$ Rankloss ($\times 10^{2}$) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
Ground-truth Training GT w/ BCE 88.90 (0.09) 5.784 (0.036) 1.649 (0.177) 94.35 (0.14) 83.70 (0.34) 87.43 (1.02) 81.91 (0.63)
IUN 82.85 (0.33) 6.311 (0.060) 2.741 (0.087) 91.21 (0.37) 71.17 (2.28) 95.04 (1.41) 60.64 (2.15)
Baseline Methods AN 65.41 (1.28) 8.403 (0.177) 10.908 (0.349) 71.74 (0.19) 27.16 (0.12) 86.42 (2.26) 16.92 (0.09)
AN-LS 64.89 (0.71) 11.175 (0.189) 21.113 (1.477) 70.81 (0.25) 20.11 (0.57) 80.53 (5.82) 12.22 (0.44)
WAN 68.70 (0.49) 7.481 (0.185) 7.575 (0.647) 78.36 (0.17) 50.14 (0.47) 84.57 (1.09) 36.95 (0.35)
EPRa 68.62 (1.87) 7.277 (0.093) 8.453 (0.390) 84.91 (0.57) 70.09 (1.12) 62.33 (0.85) 81.76 (1.78)
General Label Noise Methods ELR 69.97 (0.69) 7.176 (0.014) 5.401 (0.052) 88.25 (0.20) 59.36 (0.89) 74.95 (0.96) 56.19 (0.35)
$L$-Cp 64.71 (1.07) 7.629 (0.132) 7.426 (0.465) 78.73 (0.57) 45.64 (1.16) 76.96 (3.89) 36.35 (0.72)
$L$-Ct 67.48 (1.53) 8.302 (0.260) 6.846 (0.374) 85.78 (0.65) 55.78 (0.91) 78.65 (4.73) 48.19 (0.88)
$L$-R 69.56 (0.43) 8.314 (0.023) 7.958 (0.313) 80.75 (0.58) 51.42 (2.25) 85.28 (2.87) 39.15 (2.38)
BoostLU 68.99 (1.11) 7.218 (0.113) 5.733 (0.308) 85.79 (0.94) 54.86 (0.83) 79.39 (1.79) 47.40 (0.97)
PUL Methods Dist-PUa 71.38 (0.28) 6.493 (0.112) 3.761 (0.177) 90.02 (0.11) 60.59 (3.33) 78.60 (6.94) 57.09 (3.12)
SPML Methods ROLE 56.68 (0.79) 13.238 (0.053) 19.038 (0.294) 76.44 (0.38) 43.66 (0.64) 84.18 (3.33) 31.10 (0.49)
GR 62.69 (0.79) 11.053 (0.091) 20.911 (0.577) 72.34 (0.08) 28.62 (0.57) 83.20 (4.34) 18.36 (0.31)
LAC 62.61 (1.01) 7.746 (0.086) 9.187 (0.231) 72.74 (0.03) 28.84 (1.87) 82.41 (4.61) 18.56 (1.31)
MIME 61.96 (1.94) 7.150 (0.166) 7.016 (0.799) 73.94 (0.28) 31.96 (1.03) 81.02 (2.73) 20.82 (0.66)
Proposed AdaGC 74.82 (0.90) 6.607 (0.054) 4.381 (0.154) 86.51 (0.49) 60.68 (2.24) 77.86 (1.33) 60.03 (2.15)
@@
Table III summarizes the performance of all methods on the reBEN dataset when subjected to Random SPML noise. Our proposed AdaGC demonstrates superior overall performance, achieving the best mAP, OA, mF1, coverage, and rankloss scores. While it does not achieve the best performance in mprecision and mrecall, AdaGC shows the best balance between two, as reflected in the mF1 score.
General label noise and PUL methods show good performance in this scenario, outperforming SPML-specific methods. Additionally, WAN demonstrates to be a strong baseline, outperforming all other baselines as well as many other comparison methods. The reason might be found in the hyperparameter selection. Indeed, we tuned them for each method individually by monitoring the noisy validation mAP. While noisy mAP can be a good indicator for early learning detection when evaluated continuously during training, its fluctuations in hyperparameters selection, combined with methodologies that stem from AN-like approaches (e.g., pseudo-labelling or label correction approaches), can result in suboptimal hyperparameters’ selection. However, we argue that methods benchmarking should reflect real-world settings as much as possible, which is ensured in our experimental setting.
Comparing the results with GT training, we can observe that the AdaGC approaches the performance of training with GT multi-labels, especially in terms of mF1. Compared to IUN, AdaGC achieves a worse mAP score and a better mF1 score. Although false negatives do not affect IUN in mAP results, the mF1 scores emphasize the importance of methodologies that leverage pseudo-labels or label correction. Such approaches are vital for accurately reconstructing multi-labels, thus preventing classifier confidence bias and achieving higher mF1 scores.
@@
While Random SPML can be a good simulation for cases where both real-world single labels and GT multi-labels are difficult to gather, it provides a simplified scenario. Indeed, random single labels can still provide all the necessary information for describing the land cover, since the single labels are chosen randomly among the GT labels. On the contrary, dominant single labels are completely driven by the actual observed image, leading to cases where specific non-dominant class patterns are entirely missing from the training set, making this scenario more challenging. This can be observed in Table IV, where all methods show lower absolute performance compared to their counterparts in Table III.
Despite this increased difficulty, the proposed AdaGC maintains its leading position, achieving the highest mAP, mF1, coverage, and rankloss scores. Instead, the comparison methods struggle to outperform the WAN and EPR baseline methods. WAN shows the second best mAP score, and its mF1 score surpasses all SPML-specific methods. Only methods based on the LL family and Dist-PU are able to outperform WAN in terms of mF1, struggling against EPR. Due to the regularization term that affects the number of predicted positive labels, EPR can overcome the strong bias introduced by Dominant SPML in terms of mF1 score, despite its low mAP score.
IV-C3 AID

### v2

@@ IV-A2 AID-multilabel
It extends the AID scene classification dataset [xia_aid_2017] with multi-label annotations. The dataset comprises 3,000 high-resolution RGB aerial image patches ($600\times 600$ pixels) across 17 classes, collected from China, U.S., and several European countries. We exclude the extremely rare “mobile home” class (only 2 samples), resulting in 16 classes. Following [hua_relation_2020], 20% of the data is held out for testing. We then use 25% of the remaining training set for validation. Unlike reBEN, AID’s high spatial resolution allows multiple objects to be easily recognized, making it difficult to identify a single dominant class. Therefore, in addition to Random simulation, we manually annotated the AID dataset in a single-positive manner to assess model performance under realistic SPML conditions. For each image, four randomly shuffled candidate classes were displayed at a time to speed up annotation, with the option to refresh if none were suitable. Consecutive images were also prevented from being assigned to the same class to reduce potential annotation bias. This procedure makes the Manual labels behave similarly to random SPML labels, yet retain a slight human selection bias. The entire annotation process took approximately 4.5 hours. We also found and corrected several label inconsistencies in the original AID-multilabel data, most of which are from the meadow folder. The corrected multi-label annotations, the manual single-positive labels, and the annotation tool will be released publicly along with the AdaGC code. Dataset statistics are provided in Table II. For simplicity, in the following we use AID to refer to AID-multilabel.
@@
$\textbf{mAP}(\%)$ Coverage $\downarrow$ Rankloss (%) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
@@
RS-specific Methods RCML $64.91(0.12)$ 4.585 (0.005) 4.467 (0.020) 91.24 (0.04) 52.90 (0.32) 78.57 (1.29) 42.42 (0.34)
LCR $54.56(0.32)$ 4.864 (0.009) 5.912 (0.044) 87.16 (0.03) 28.52 (0.14) 73.82 (1.81) 19.31 (0.08)
Proposed AdaGC $67.85(0.05)$ 4.553 (0.012) 4.165 (0.033) 92.49 (0.01) 64.46 (0.24) 70.43 (0.59) 63.22 (0.10)
@@
IV-B Experimental Setup
We adopt ResNet-34 and ResNet-50 as backbones for the reBEN and AID datasets, with batch sizes of 512 and 32, respectively. ResNet-34 is chosen for reBEN as an efficient compromise, since our focus is on evaluating method effectiveness rather than maximizing absolute accuracy. For AID, we use ImageNet-pretrained weights given its RGB input. To adhere to a realistic weakly supervised scenario where ground-truth multi-labels are unavailable, we strictly avoid the use of clean data for model selection. Consequently, all hyperparameters are tuned via grid search on the noisy validation set, simulating the absence of clean labels. While this strategy relies on imperfect supervision, it prevents the introduction of “oracle” information and ensures that the reported performance reflects the method’s capability to generalize without privileged access to clean labels. Despite the use of noisy validation labels, the resulting metrics remain informative for model selection. Under SPML noise, random errors are inherently difficult for deep models to memorize, and thus the validation signal largely reflects the underlying clean structure. As confirmed in our experiments, noisy mAP validation scores maintain a strong correlation with clean test performance, making them a practical and reliable indicator in the absence of ground-truth labels. To reduce computational cost, tuning is performed on subsets (20% for reBEN, 50% for AID) and fewer epochs (10) for comparison methods. For AdaGC, the Mixup Beta parameter $\alpha$ is set to 1. The GC loss weight $\lambda$ is set to 3, 5, 9, and 8 for reBEN-Random, reBEN-Dominant, AID-Random and AID-Manual, respectively, and the learning rate is set to 1e-3 and 5e-4 for reBEN and AID, respectively, based on tuning results.
For the final evaluation, all models are trained for 70 epochs on the full training sets using the optimal hyperparameters and evaluated on the clean test sets. To ensure robustness and statistical reliability, each experiment is repeated for three Monte Carlo runs. We report the mean and unbiased standard deviation of all metrics.
@@ IV-B Experimental Setup
$\textbf{mAP}(\%)$ Coverage $\downarrow$ Rankloss (%) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
@@
RS-specific Methods RCML $58.26(0.82)$ 5.890 (0.134) 8.035 (1.181) 88.73 (1.05) 40.75 (1.27) 83.06 (2.35) 30.93 (2.88)
LCR $55.64(0.20)$ 6.579 (0.045) 8.151 (0.080) 88.67 (0.00) 34.25 (0.06) 83.24 (0.54) 24.05 (0.04)
Proposed AdaGC $60.06(0.16)$ 5.581 (0.026) 6.877 (0.054) 90.00 (0.06) 59.08 (0.25) 60.19 (0.47) 63.43 (0.37)
@@
In addition to mAP, we report two ranking-based metrics, i.e., coverage and ranking loss (Rankloss), and four accuracy-based metrics, i.e., overall accuracy (OA), mean F1 score (mF1), mean precision (mprecision), and mean recall (mrecall). The metrics are defined as follows.
  • mF1 is the average of F1-scores computed per class:
  • Baseline methods: We considered several baselines [Cole_2021_CVPR], such as AN, AN-LS, Weak (WAN), and EPR. All use a BCE loss at their core, while EPR uses only the BCE positive term and adds a regularization term, assuming the a-priori knowledge of the average number of positive labels per sample (see Tables I and II).
  • General label noise methods: From this category, we considered the ELR approach adapted as in [Burgert_2022_TGRS], the LL family of approaches (i.e., LL-R, LL-Ct, and LL-Cp) [Kim_2022_CVPR], and the BoostLU [Kim_2023_CVPR] in combination with LL-Ct.
  • PUL methods: We considered a single state-of-the-art (SOTA) method from this category, i.e., Dist-PU [Zhao_2022_CVPR], adapted for SPML by simply applying it to each class independently.
  • SPML-specific methods: The main comparison methods we considered for SPML are ROLE [Cole_2021_CVPR], GR Loss [Chen_2024_IJCAI], LAGC [Xie_2022_NIPS], and Mutual label enhancement for sIngle-positive Multi-label lEarning (MIME) [Liu_2023_ICML].
  • Upper bound methods: For analytical comparison with GT training, we considered two cases: i) training with GT multi-labels using a BCE loss (GT w/ BCE), and ii) the with true Negatives (IUN) approach [Cole_2021_CVPR], where in addition to the single positive labels, all true negatives are provided for training. IUN allows us to assess the performance in a hypothetical scenario where all true negatives have been retrieved. False negatives are ignored during training, meaning that the model is still provided with a single positive label for each image.
  • RS-specific methods: We also included RCML [Aksoy_2024_TNNLS] and LCR [hua_learning_2020], which were originally developed for general noisy multi-label RS classification. RCML leverages a collaborative learning framework with a ranking loss for sample selection to mitigate label noise, while LCR uses word embedding similarity to regularize model training under noisy annotations. To adapt them to SPML, we use only the missing-class labels term in RCML and perform selection per class rather than per sample. For LCR, we use RemoteCLIP’s [liu_remoteclip_2024] text encoder to encode multi-word class names. These adaptations ensure a fair comparison while preserving the core noise-robust strategies of the original methods.
IV-C Benchmark Results
This section analyzes the benchmark results reported in Tables IIIVI. Across all scenarios, AdaGC consistently achieves the highest mAP and strong SPML performance in terms of coverage, rankloss, and mF1, demonstrating its effectiveness and robustness in handling SPML in RS imagery. The upper-bound IUN performs well in mAP but suffers from low mrecall due to its inherent single-positive bias. AN and AN-LS are highly sensitive to false negatives, exhibiting high mprecision but low mrecall and consequently poor mAP. In contrast, WAN is more robust and achieves the best baseline mAP by better balancing precision-recall trade-offs. EPR, trained exclusively on positive labels with a cardinality regularizer, consistently shows the highest mrecall across datasets but low mprecision, reflecting its tendency to overestimate positives. Its advantage largely stems from the use of the true average number of labels per image, a prior not shared by other competitors. The two RS-specific methods designed for general label noise exhibit clear limitations under SPML conditions. The following subsections detail dataset- and scenario-specific observations.
@@ IV-C Benchmark Results
TABLE V: Test Performance Comparison of Different Methods on the AID-Random Dataset. The Best Average Metric Values Are Reported in Bold, Second Bests Are in Italic. The Related Standard Deviations Are Reported in Brackets
$\textbf{mAP}(\%)$ Coverage $\downarrow$ Rankloss (%) $\downarrow$ OA (%) $\uparrow$ mF1 (%) $\uparrow$ mprecision (%) $\uparrow$ mrecall (%) $\uparrow$
Ground-truth Training GT w/ BCE $89.01(0.32)$ 5.834 (0.011) 1.657 (0.061) 94.29 (0.19) 83.12 (0.66) 88.87 (0.38) 79.58 (1.28)
IUN $82.24(0.61)$ 6.622 (0.031) 3.408 (0.026) 90.88 (0.19) 72.65 (0.33) 93.13 (0.86) 62.87 (0.49)
Baseline Methods AN $64.89(1.28)$ 8.557 (0.164) 11.360 (0.351) 71.60 (0.23) 26.83 (0.18) 86.05 (2.18) 16.71 (0.13)
AN-LS $64.40(0.73)$ 11.347 (0.231) 21.664 (1.657) 70.65 (0.28) 19.97 (0.63) 79.98 (5.71) 12.14 (0.47)
WAN $68.12(0.46)$ 7.620 (0.173) 8.055 (0.588) 78.06 (0.16) 49.53 (0.42) 84.19 (1.05) 36.45 (0.27)
EPRa $68.38(1.84)$ 7.395 (0.079) 9.086 (0.412) 84.76 (0.57) 69.82 (1.10) 62.30 (0.85) 81.11 (1.68)
General Label Noise Methods ELR $69.13(0.73)$ 7.469 (0.063) 5.914 (0.152) 87.99 (0.12) 59.04 (0.84) 74.80 (1.00) 55.91 (0.30)
$L$-Cp $64.42(1.06)$ 7.792 (0.141) 7.931 (0.471) 78.64 (0.58) 45.47 (1.25) 76.83 (3.94) 36.25 (0.78)
$L$-Ct $66.96(1.60)$ 8.574 (0.243) 7.647 (0.322) 85.42 (0.65) 55.41 (0.91) 78.43 (4.80) 47.92 (0.75)
$L$-R $68.85(0.44)$ 8.526 (0.018) 8.616 (0.294) 80.39 (0.58) 50.89 (2.19) 84.92 (2.85) 38.75 (2.31)
BoostLU $68.00(0.97)$ 8.104 (0.127) 6.656 (0.519) 85.44 (0.97) 54.48 (0.86) 79.19 (1.90) 47.14 (1.00)
PUL Methods Dist-PUa $70.85(0.47)$ 6.692 (0.133) 4.355 (0.265) 89.86 (0.09) 60.48 (3.26) 78.55 (6.93) 57.04 (3.10)
SPML Methods ROLE $56.35(0.77)$ 13.581 (0.041) 19.800 (0.282) 76.10 (0.38) 43.26 (0.73) 83.73 (3.32) 30.84 (0.57)
GR $62.14(0.73)$ 11.236 (0.090) 21.537 (0.567) 72.11 (0.09) 28.32 (0.56) 82.59 (4.41) 18.18 (0.32)
LAC $62.27(0.92)$ 7.902 (0.091) 9.674 (0.266) 72.57 (0.01) 28.59 (1.81) 82.17 (4.10) 18.38 (1.22)
MIME $61.55(1.86)$ 7.308 (0.175) 7.474 (0.756) 73.67 (0.26) 31.55 (1.32) 81.05 (3.42) 20.52 (0.86)
RS-specific Methods RCML $64.24(1.41)$ 7.522 (0.107) 8.338 (0.445) 80.10 (0.64) 45.45 (0.47) 75.62 (4.66) 37.84 (0.23)
LCR $62.62(0.47)$ 8.269 (0.165) 10.453 (0.500) 71.68 (0.13) 23.48 (1.01) 78.12 (2.50) 15.06 (0.58)
Proposed AdaGC $75.67(0.26)$ 6.862 (0.209) 5.636 (1.303) 83.39 (2.40) 64.16 (1.57) 68.89 (5.58) 71.82 (6.80)
@@
Table III presents the results on the reBEN dataset under Random SPML noise. AdaGC delivers the strongest overall performance, achieving the highest mAP, OA, and mF1, together with the best coverage and rankloss scores. Although not optimal in mprecision or mrecall individually, it provides the most effective precision–recall trade-off, as reflected by its mF1. General label-noise and PUL methods perform well in this setting, often surpassing SPML-specific approaches. WAN remains particularly competitive among baselines. Within RS-specific methods, RCML offers noticeable improvements compared to baselines, whereas LCR contributes little due to its fixed class-correlation assumptions. Compared with GT training, AdaGC closely approaches the full-label upper bound, especially in mF1. Compared to IUN, AdaGC attains slightly lower mAP but higher mF1. This highlights the importance of incorporating pseudo-labeling or label-correction mechanisms to recover missing labels and avoid confirmation bias.
@@
Random SPML provides a simplified single-positive noise setting, whereas Dominant SPML is a more practical scenario driven by the visually dominant class in each image, resulting in systematic omission of non-dominant classes. This significantly increases task difficulty, as reflected in Table IV, where all methods exhibit lower absolute performance compared with Table III. Despite this challenge, AdaGC remains the strongest method, achieving the best mAP, mF1, coverage, and rankloss. In contrast, most competing approaches fail to surpass the WAN and EPR baselines. WAN obtains the second-best mAP and outperforms all SPML-specific methods in mF1. LL-based methods and Dist-PU offer additional gains over WAN in mF1, yet still fall short of EPR. Thanks to its cardinality regularization, EPR effectively mitigates the strong bias induced by Dominant SPML, yielding competitive mF1 despite its low mAP. RCML remains relatively strong but continues to underperform the proposed AdaGC.
IV-C3 AID-Random